Carswell Electronics adopted the dollar-value LIFO method on January 1, 2021, when the inventory value of its one inventory pool was $870,000. The company decided to use an external index, the Consumer Price Index (CPI), to adjust for changes in the cost level. On January 1, 2021, the CPI was 230. On December 31, 2021, inventory valued at year-end cost was $1,045,000 and the CPI was 253.
 
Required:
Calculate the inventory value at the end of 2021 using the dollar-value LIFO method. (Round "Year end cost index" to 2 decimal places.)
Answer #1
date

inventory at

 year-

end

 cost

/

year-

end 

cost

 index

=

inventory 

layers at

 base 

year cost


inventory 

layers at

 base year 

cost

x

year-end 

cost 

index

=

inventory 

layers

 converted to

 cost

ending inventory DVL cost
1/1/21870,000/1=870,000base870,000x1=870,000870,000













12

/31

/21

1,045,000/1.1=950,000base870,000x1=870,000






202180,000x1.1=88,000958,000

253/230= 1.1

950,000-870,000= 80,000

answered by: anonymous